Find the value of x,
x+5+3x+6x = 50

Found 2 solutions by stanbon, Cromlix:
Answer by stanbon(75887)

Find the value of x,
x+5+3x+6x = 50
--------
10x + 5 = 50
10x = 45
x = 4.5
------------
Cheers,
Stan H.
--------

Answer by Cromlix(4381)

Hi there,
x+5+3x+6x = 50
Collect like terms
x + 3x + 6x = 50 - 5
10x = 45
x = 45/10
x = 4.5
Hope this helps:-)